OpenStudy (goldphenoix):

Lance is 39 years old. Ted is 22 years old. Rosalinda is 45 years old. How old is Erik if the mean age of the four friends is 33.5? (I forgot how to solve these problems >_>)

OpenStudy (anonymous):

mean age means average

OpenStudy (anonymous):

so that means lance's age + ted's age + rosalinda's age + erik's age all divided by 4 = 33.5

OpenStudy (anonymous):

do you have an idea how to get erik's age now?

OpenStudy (anonymous):

\[33.5 = \frac{ 39 + 22 + 45 + x }{ 4 }\]

OpenStudy (goldphenoix):

Oh yes. Now it looks familiar. Add 39 + 22 + 45 to get 106 Now it should look like 33.5 = 106 + x / 4?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

but 106 + x should be a quantity

OpenStudy (goldphenoix):

Then you do 33.5 * 4 to get 134. Now it should look like: 134 = 106 + x ?

OpenStudy (anonymous):

:)

OpenStudy (goldphenoix):

I'm guessing smiley face means yes. Now subtract 134 from 106 to get 28. Equation should look like 28 = x So that means that Erik age is 28, yes?

OpenStudy (anonymous):

yep you can even check your work by finding the mean

OpenStudy (goldphenoix):

39 + 22 + 45 + 28 = 134 134 / 4 = 33.5 Alright, 28 is the answer. Thanks for your help, hobbs and xlegendx. I'm gonna give the medal to xlegendx, since you already have a medal, hobbs. ;)

OpenStudy (anonymous):

all good as long as you learned <3